For each of the folla ing situations, give the degrees cr ee m the group DFG), f error DFE), and f the total DFT State the null and alternative hypotheses, H and Ma and give the numerator and de om i at de ees off eedo f the F atistic. () A poultry farmer is interested in reducing the cholesterol level in his marketable eggs. He wants to compare two dTerent cholesterol-lowering drugs edded to the hens stendard diet es well as an all-vegetarien diet. He assigns 25 of his hers to eech of the three trestments DFG- OFT Ho: The al vegetarian diet group has a ower mean cholestarol level o At leest one group hes a different mean chalesterol level o The al-vegctarian diet group has a higher mean cholesterol Ievel O All groups have the same mean cholesterol lavel. o All groups heve different mean cholesterol levels. a: The al egetarian diet group has a higher mean cholestorol leval. O All groups have the same mean chalesterel lvel At least one group has different mean cholesterol level. o All groups have cifferent mean cholesterol levels o The all vegetarian diet group has a ower mean cholestarol level. numerator d b) A researcher is nterested in students opinions regarding an additional annual ce to support non-income-p duong varsity soorts. Students were asked to ate ther อ0ceptance o this ee on attend varsity football or basketball games only, 16 were from students who also attend other varsity competitions, and 48 who did not attend any varsity games. DFG OFE OFT seven-point scale She received 94 esponses of which 30 were rom stu s who Ho: O All groups have cifferent mean ratings O All groups heve the same mean rating O The group of students who do not attend games has ง higher mean rating. The group af stuCents who do not attend games has a lower mean rating. o At least one group hes a different mean rating Ma: o The group af stuCents who do not attend games has a lower mean rating. At least one group has a cifferent mean rating. o The group of students who do not attend games has a higher mean rating. Al groups have the same mean rating. A groups have cirferent mean ratings.
Answer #1

a)

DFG =2

DFE =25*3-3 =72

DFT =74

Ho: all group have the same

Ha:at least one group has a different mean

numerator df =2

denominator df =72

b)

DFG =2

DFE =94-3 =91

DFT =93

Ho: all group have the same

Ha:at least one group has a different mean

numerator df =2

denominator df =91